OpenStudy (anonymous):

\(\text{For an odd prime p, verify that the sum}\\ \sum _{i=1}^{(p-1)}i^n= 0 \mod p \text{ or } \sum _{i=1}^{(p-1)}i^n= -1 \mod p\)

ganeshie8 (ganeshie8):

*

OpenStudy (anonymous):

xD

OpenStudy (jhannybean):

*

OpenStudy (anonymous):

-.-

Parth (parthkohli):

*

OpenStudy (anonymous):

now @Kainui and @PFEH.1999 left lol if they click on star as well no one left to solve

Parth (parthkohli):

Check for \(n=1\).\[\sum _{i = 0}^{p-1}i = \frac{p(p-1)}{2} \equiv 0 \pmod{p}\]Maybe induction?

OpenStudy (anonymous):

why not try if it works :D

ganeshie8 (ganeshie8):

thats a good idea !

ganeshie8 (ganeshie8):

since p is odd, 2|(p-1) consequently the sum is a multiple of p

Parth (parthkohli):

Yup, but that is only \(n=1\).

ganeshie8 (ganeshie8):

that clears the base case

OpenStudy (anonymous):

well trying to know which numbers divisable by p-1 is a good step

ganeshie8 (ganeshie8):

the trick part is remainder is not 0 for all \(n\)

OpenStudy (anonymous):

hmm its not hard though if u tried

ganeshie8 (ganeshie8):

it is easy if you allow me to use fermat's little thm

OpenStudy (anonymous):

i think its solvable with fermat :D

ganeshie8 (ganeshie8):

im still trying the induction idea though as it looks interesting

ganeshie8 (ganeshie8):

By Fermat's little thm, \(a^{p-1}\equiv 1\pmod {p}\) it follows \(a^n \equiv 1 \pmod {p}\) whenever \((p-1)|n\) so the remainder for the sum in question would be \(1 + 1+ \cdots + (\text{p-1 times}) = p-1\equiv -1 \pmod{p}\)

ganeshie8 (ganeshie8):

need to work the case when \((p-1) \not |~ n \)

OpenStudy (anonymous):

yaaay

OpenStudy (mathmath333):

\(\)

Parth (parthkohli):

You spelled "*" wrong.

Parth (parthkohli):

If we're still at the induction thingy, http://en.wikipedia.org/wiki/Faulhaber%27s_formula

Parth (parthkohli):

I'm trying to see if that formula tells us anything about divisibility.

OpenStudy (anonymous):

yeah u revealed zeta :P but its ok , u can do that as well i thought using NT would be much easy to make it introduction to zeta .

OpenStudy (mathmath333):

what is the maximum value for \(n\)

OpenStudy (anonymous):

n can be anything

Parth (parthkohli):

OK, yes, it seems that there's an (n+1) term in every expansion. So our sum is divisible by p?

OpenStudy (anonymous):

ganesh already proved case one with p-1|n it gives -1 remainder

OpenStudy (anonymous):

u can mix with binomial idk there is many ways to prove its zero when p-1 dont divide n pk try this by induction

OpenStudy (anonymous):

well i checked induction works with first case :D

OpenStudy (mathmath333):

for case 1 wolfram is giving false https://www.wolframalpha.com/input/?i=%5Csum+_%7Bi%3D1%7D%5E%7B%285-1%29%7Di%5E%7B100%7D+%28mod+5%29%3D+0+

OpenStudy (anonymous):

assume p-1|n then p-1 dont divide n+1 (p-1) even so we got by fermat (1+2+3+4+.....+p-1)=p(p-1)/2 mod p=0 mod p since p-1 is even now we need to try up to n+p

OpenStudy (anonymous):

@mathmath333 case 1 when p-1|n it gives -1 mod not 0 :D

ganeshie8 (ganeshie8):

when \((p-1)\not | ~n\) : Notice the bases in the sum \(B = \{1,2,3,\ldots,(p-1)\}\) are relatively prime to \(p\) Next consider an integer \(r\) which is relatively prime to \(p\) and whose order is \(p-1\) Since \(r\) is relatively prime to \(p\), the same holds for all powers of \(r\). Hence \(r^k\) is congruent to some positive integer \(\in B\) so the sum \(r^{0n}+r^{1n} + r^{2n} + \cdots + r^{n(p-2)}\) will be congruent to the sum in question. \[r^{0n}+r^{1n} + r^{2n} + \cdots + r^{n(p-2)} = \dfrac{1-r^{(p-1)n}}{1-r^n} \equiv 0 \pmod {p}\] \(\blacksquare\)

OpenStudy (anonymous):

how did u get last step ?

OpenStudy (anonymous):

yes yes

ganeshie8 (ganeshie8):

\[ \dfrac{1-r^{(p-1)n}}{1-r^n} \] do you agree the nujmerator is divisible by \(p\) ?

OpenStudy (anonymous):

its perfect solution though :P yes

OpenStudy (anonymous):

i thought u used geometric series in like inverse way maybe xD

ganeshie8 (ganeshie8):

and the denominator is not divisible by \(p\)because we assumed that \((p-1) \not | ~n\) consequently the fraction is divisible by \(p\)

OpenStudy (anonymous):

ik ik how did u establish that \(r^{0n}+r^{1n} + r^{2n} + \cdots + r^{n(p-2)} = \dfrac{1-r^{(p-1)n}}{1-r^n} \)

ganeshie8 (ganeshie8):

it helps to think of the fraction like this : \[\dfrac{1-r^{(p-1)n}}{1-r^n} = \dfrac{(1-r^n)\times p\times (stuff)}{1-r^n}\]

OpenStudy (anonymous):

hehe yeah amazing :D

OpenStudy (anonymous):

remember this approach , will help incase u decided to learn zeta function ;)

ganeshie8 (ganeshie8):

i skipped many steps in my reply to keep it short, il be glad to try to answer if anyone has questions.. :)

ganeshie8 (ganeshie8):

teach me zeta function @Marki

ganeshie8 (ganeshie8):

maybe in a new post..

OpenStudy (mathmath333):

so when \(n-1| p\) it is \(-1\) and when \(n-1\not | p\) it is \(0\) ?

ganeshie8 (ganeshie8):

Exactly ^ still trying for a counter example ha :P

OpenStudy (mathmath333):

haha

OpenStudy (anonymous):

haha wont find one its solid theorem :D

ganeshie8 (ganeshie8):

when \((p-1)|n\), the remainder is \(-1\) otherwise it is \(0\)

OpenStudy (anonymous):

yes yes , btw induction works good , start from p-1 |n then check for z_(p-1) n+1 n+2 n+3 n+4 ....... up to n+p

OpenStudy (anonymous):

at each time u'll gonna find geometric series which is known it divide p

OpenStudy (anonymous):

some way :P

OpenStudy (anonymous):

but i found @ganeshie8 solution more sense

OpenStudy (anonymous):

:D thats it

ganeshie8 (ganeshie8):

induction looks neat, but i missed few replies above while cooking up the solution.. let me scroll up read them again :)

ganeshie8 (ganeshie8):

so you have induction proof for both cases ?

OpenStudy (anonymous):

well u agree that DA gives all states?

OpenStudy (anonymous):

so first case is p-1|n gives -1 from Fermat .. then we need to show by induction that other cases divides

OpenStudy (anonymous):

or we need to show it gives 0 r -1 (ovs all gives 0)

ganeshie8 (ganeshie8):

got it ! you're inducting on n, pretty neat :)

OpenStudy (kainui):

*

OpenStudy (anonymous):

lol kai its already been solved :P

OpenStudy (kainui):

I was just joking around =P